answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

A science class 12 boys and 15 girls What is the probability that a student chosen at random is a boy?

12/27 reduces to 4/9

numerator = the number of acceptable outcomes

denominator = total number of outcomes (12 boys + 15 girls)

What is the probability that in a group of six people at least two were born in the same month?

For ease of answering, we will work under the assumption that the probability of someone being born within any given month is equal to that of any other month.

Allowing that assumption, we can look at that question a slightly different way and say "What is the probability that all people in a group of six would each be born in a different month?" The answer to that would be 12/12 * 11/12 * 10/12 * 9/12 * 8/12 * 7/12, which can also be expressed as (12! / 6!) / 126, and comes out to 665280 / 2985984, which equals 385 / 1728.

The probability of at least two being being born in the same month would then be:

1 - 385 / 1728

= 1343 / 1728

≈ 0.7772, or approximately 77.72%

What is the probability of getting two heads and one tail when a fair coin is tossed three times?

If the sequence matters (you want H-H-T in that order), then . . .

Probability of the first head = 0.5

Probability of the second head = 0.5

Probability of the tail on the 3rd toss = 0.5

Probability of the correct 3-toss sequence = (0.5 x 0.5 x 0.5) = 1/8 = 0.125 = 12.5%

=====================================================

If the sequence doesn't matter, then the probability is higher.

All the possible results of 3 tosses are:

TTT

TTH

THT

THH

HTT

HTH

HHT

HHH

If the sequence doesn't matter, there are 3 different ways to get 2 heads and 1 tail.

The probability is

3/8 = 0.375 = 37.5%

What is the probability that the sequence will be head-head-tail when a fair coin is tossed three times?

Probability of the first head = 0.5

Probability of the second head = 0.5

Probability of the tail on the 3rd toss = 0.5

Probability of the correct 3-toss sequence = (0.5 x 0.5 x 0.5) = 1/8 = 0.125 = 12.5%

What is the probability of getting either a sum of 6 or atleast one 6 in the roll of pair of dice?

Each cube can fall in 6 different ways.

Two cubes can fall in (6 x 6) = 36 ways.

16 different successes:

1, 5

2, 4

3, 3

4, 2

5, 1

1, 6

2, 6

3, 6

4, 6

5, 6

6, 6

6, 1

6, 2

6, 3

6, 4

6, 5

Probability of success = 16/36 = 44.4%(rounded)

In research how to define standard deviation?

Standard deviation shows how much variation there is from the "average" (mean). A low standard deviation indicates that the data points tend to be very close to the mean, whereas high standard deviation indicates that the data are spread out over a large range of values.

A multiple choice quiz consists of 6 questions each with 4 possible answers If a student guesses at the answer to each question then the mean number of correct answers is?

If there are four possible answers to a question, then a guessed answer would have a probability of 1 in 4. If there are six questions, then the mean number of correct answers would be six times 1 in 4, or 1.5

Why is the number twenty three everywhere?

Many interesting coincidences have been observed, relating to the number 23 (the late Robert Anton Wilson wrote about this at some length, and of course there is even a movie about it) but there are also lots of interesting coincidences concerning many other numbers, and chances are, something notable can be observed about practically any number you come up with. We are always counting things and attaching numbers to things, everything happens at a specific time that can be described by a number, and a specific place that has an address, and involves a specific number of people and objects, having specific weights, speeds, monetary worth, and various other factors all of which can be described by a number; there are endless opportunities to find numbers and to find coincidental relationships between numbers. Consider the fact that in the radio comedy program (later made into novels, a TV series, and a movie) called "The Hitchhiker's Guide To The Galaxy, Douglas Adams suggested that the ultimate answer to life, the universe, and everything is 42. This was a joke, however, many people heard or read this, and started looking more closely at the number 42, and guess what, there are all sorts of interesting coincidences involving the number 42. It works just as well as the number 23. And it was chosen at random.

What does it means if the standard deviation is small?

Simplest terms, it means that the sample values are grouped closely together. Another way of saying it is that the range of values is small.

The plot will be a leptokurtic plot (peaked).

A bag contains 8 red marbles 6 blue marbles and 9 green marbles If a marble is selected at random what is the probability of choosing a blue marble?

Number of different possible choices = 8 + 6 + 9 = 23

Number of available successful choices (blue marbles) = 6

Probability of success = 6/23 = 0.26087 = 26.087 %(rounded)

What is the probability of getting tails on all 7 tosses if a fair coin is tossed 7 times?

The probability of Tails on the first toss is 1/2.

The probability of Tails on the second toss is 1/2.

The probability of Tails on the third toss is 1/2.

The probability of Tails on the fourth toss is 1/2.

The probability of Tails on the fifth toss is 1/2.

The probability of Tails on the sixth toss is 1/2.

The probability of Tails on the seventh toss is 1/2.

The probability of all of them is (0.5 x 0.5 x 0.5 x 0.5 x 0.5 x 0.5 x 0.5) = (0.5)7 = 0.0078125

= 0.78125 %

Distinguish between a gross sample and a laboratory sample?

In geology say, a gross sample would be a representative sample of the whole rock. A lab sample would be a fraction of that, refined such that measurements are able to be made on a single mineral.

for example, a piece of basalt would be a convenient field (gross) sample, from which say, mica is extracted after grinding and separation, to enable the 'date' of the basalt to be determined.

Similar concepts would apply in other fields, such as biology, botany, water sampling and so on.

You spin a spinner two times find the probability that the spinner stops on 3 then 1?

The probability that a spinner with N sides stops on 2 particular numbers in two spins in 1 in N2. It does not matter what the two numbers are, since the two spins are sequentially unrelated.

Trending Questions
In probability is heads and tails the same as tails and heads? What is Mutually Exclusive probability? You roll 2 dice You can use a Venn diagram to calculate the probability that the sum of the dice is a prime number? When you ask someone for a number between one and ten why do they always say seven? A certain airplane has two independent alternators to provide electrical power The probability that a given alternator will fail on a 1-hour flight is 02 What is the probability that a both will fail? Is a parallelogram's diagonals perpendicular? What does it mean when a man introduce his daughter to a woman? How many different 5 cards hands would contain exactly one ace? The number in which the letters in the word party can be arranged? What is the probability that the sum of the numbers from two rolled dice result in a perfect square or even number? What is probability of exceedance? How many ways can you arrange the letters of the word NUMBERS? You have 6 vertices and 5 faces? What are the odds of rolling a four on one die? What is the chances of winning the lottery? Is 0.006 a probability? Which is more consistency arthematice mean is 110 and standard deviation is 25 and arthematic mean is 90 and standard deviation is 15? If 10 percent of all disk drives produced on an assembly line are defective what is the probability that there will be exactly one defect in a random sample of 5 of these? Two dice are rolled 9 times with a total sum of sixty. What is the mean sum? What is the probability of rolling a 4 twice with two rolls of a number cube?